Gould Corporation uses the following activity rates from its activity-based costing to assign overhead costs to products:

Activity Cost Pool Activity Rate
Setting up batches $ 60.06 per batch
Processing customer orders $ 73.26 per customer order
Assembling products $ 4.75 per assembly hour

Data concerning two products appear below:

Product K91B Product F65O
Number of batches 99 70
Number of customer orders 49 63
Number of assembly hours 503 910

How much overhead cost would be assigned to Product K91B using the activity-based costing system? (Round your intermediate calculations and final answers to 2 decimal places.)

Multiple Choice

  • $11,924.93

  • $5,945.94

  • $138.07

  • $107,324.37
Answer #1
$11,924.93
Dear Student
Thank you for using homeworklib
Please find below the answer
Statement showing Computations
Particulars Product K91B
Overhead cost:
Setting up batches = 60.06*99            5,945.94
Processing customer orders = 73.26*49            3,589.74
Assembling products = 4.75*503            2,389.25
Overhead cost assigned to Product K91B          11,924.93